OpenStudy (anonymous):

4. There is a circular garden in the middle of a square yard. The radius of the circle is 4x. The side length of the yard is 20x. What is the area of the part of the yard that is not covered by the circle?

OpenStudy (callisto):

area of circle = π(4x)^2 = (16x^2)π area of square = (20x)^2 = 400x^2 area required = 400x^2 - (16x^2)π =work it out yourself!

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome